Your manager is proposing a new investment. If the company pursues the investment, it will cost $300,000 today. It has an expected life of 10 years and no salvage value. An environmental impact fee of $25,000 must be paid to the government every five years, once at the end of year five and again to safely dispose of the investment at the end of year 10. As a result of the investment, annual revenue increases are estimated to be $70,000 per year. Your company must borrow 1/2 of the initial investment cost. The bank has agreed to four equal annual payments, with the first payment due at the end of year 1. The company's MARR is 10% compounded annually. The loan interest rate is 12% compounded annually.

a) How much is the loan payment?

b) What is the present worth of this investment opportunity?

c) Based on the present worth analysis, should the company make the investment? Why or why not?
Answer #1

Ans.

The Initial Investment: $ 300000

Borrowed Amount: $ 150000 (Half of the investment)

(a) Loan Interest Rate: 12%

The Present Value Annuity Factor at 12% for 4 years is 3.0373

Therefore, Present Value= Annuity * Present Value annuity factor

                 150000 = A * 3.0373

                  A = 150000/3.0373

                  A = $ 49385.96

Therefore installment for per year loan repayment will be $ 49385.96. In total company needs to pay 4* 49385.96 = $ 197543.87.

(b) Present Worth of the investment opportunity = Aggregate of present value of cash inflow - Aggregate of present value of cash outflow

Types of cash flow Year

Amount  

(in Dollar)    (A)

PV Factor @ 10%

(B)

Present      Value

(A*B)

Cash Inflow From Year 1 to 10 (Annuity) 70000 6.1446 430122
Cash Outflow Year 0 (Initial Investment at present) 300000 1 (300000)

Cash Outflow

(Environmental Impact Fee)

at the end of year 5 25000 .6209 (15523.03)
Cash Outflow at the end of year 10 25000 .3855 (9638.58)

Present Worth of investment opportunity = 430122-(300000+15523.03+9638.58) = $ 104960.4

(c) Since the Present worth of investment opportunity is positive, therefore the project seems to be profitable. Also the interest amount on loan is lower than the present worth,therefore, it is a profit making project even after borrowing half of the investment amount.